# Electromechanical computational model of the human stomach

### v1

@@ Abstract
[5]organization=Department of Engineering, Campus Bio-Medico University of Rom, addressline=Via A. del Portillo 21, city=Rome, postcode=00128, country=Italy
@@ 1 Introduction
The stomach plays a central role in the gastrointestinal tract. It is concerned with storing, mixing, and grinding food. Fundamental to this functionality are rhythmic gastric contractions, which generate peristaltic waves that mix and propel food through the digestive tract. Perturbations of this system are associated with a broad spectrum of diseases and disorders, including obesity, which itself is a major risk factor for conditions such as cardiovascular disease. Other gastric motility disorders include gastroparesis, gastroesophageal reflux disease (GERD), and functional dyspepsia, all of which are linked to abnormal gastric motility [1, 2]. A detailed understanding of gastric motility is therefore essential for the effective diagnosis and treatment.
@@
Anatomically, the stomach is divided into three main regions (fundus, corpus, antrum) and exhibits both a so-called greater and lesser curvature as illustrated in Figure 1. Despite its relatively thin wall (on average $2.84.5\text{\,}\frac{\mathrm{mm}}{}$ in humans [3, 4]) compared to its overall size (approximately $1525\text{\,}\frac{\mathrm{cm}}{}$ in length [3, 4]), the gastric wall consists of distinct layers. The muscular layer, primarily responsible for gastric contractions, comprises two smooth muscle sheets: one oriented circumferentially and the other longitudinally. An additional oblique smooth muscle layer is restricted to the upper corpus near the esophagus. It is neglected in this study, as it plays a limited role in gastric peristalsis, which occurs mainly in the lower body and antrum [5].
@@
The stomach exhibits intricate patterns of mechanical contraction [5]. Gastric motility is governed by complex electromechanical couplings comprising at least two cell types, interstitial cells of Cajal (ICCs) and smooth muscle cells (SMCs) [6]. ICCs, arranged in a distributed network, act as pacemaker cells generating electrical activity in the form of slow waves [7]. Being self-excitatory, ICCs show peculiar features of nonlinear coupled oscillators [8, 9]; i.e., they generate electrical activity at specific intrinsic frequencies while the integrated network synchronizes at a unified emerging frequency. This so-called entrainment generates ordered waves at the highest intrinsic frequency of approximately $3\text{\,}\frac{\mathrm{cpm}}{}$ (cycles per minute). The highest intrinsic frequency is observed near the pacemaker region at the mid-corpus along the greater curvature. The intrinsic frequency decreases longitudinally and circumferentially and assumes its minimum near the pylorus at the lesser curvature [2]. The SMCs are electrically coupled to ICCs, such that the electrical slow waves coordinate the contractions of the SMCs via an elaborate excitation–contraction mechanism [10, 11]. These contractions can produce substantial gastric wall deformations, with amplitudes measured in the range of $720\text{\,}\frac{\mathrm{mm}}{}$ [12, 13, 14].
@@
Early whole-organ modeling approaches of the stomach focused primarily on its electrophysiological system [15]. More recently, models incorporating mechanical phenomena have also been proposed. Our previous work [16] was the first to simulate gastric electromechanics on an idealized whole-organ geometry, reproducing key phenomena. However, it relied on a highly simplified geometry and material model of the tissue and could deal only with very limited contraction amplitudes due to a limited numerical robustness that resulted from the use of membrane finite elements to discretize the thin wall of the stomach. [17] investigated a realistic stomach geometry. However, their model was limited to passive mechanical behavior. [18, 19] pursued a similar purely passive approach to study patient-specific bariatric surgery. [20, 21] studied contraction waves governed by an electro-chemomechanical model in the porcine stomach, using an active-stress formulation and incorporating mechano-electric feedback. While their approach was particularly advanced in terms of the electro-chemical components, several limitations remained such as an unrealistic antrum activation and mild contraction intensity.
@@
The thin-walled structure of the stomach lends itself naturally to reduced-dimensional formulations, such as membrane or shell models. While membrane models are computationally efficient, they neglect bending stiffness, limiting their application in regions with high curvature and large contractions. Nonlinear shell formulations capture both membrane and bending effects under large deformations and are thus preferable for accurate gastric biomechanics. Shell models have proven effective for studying the biomechanics of thin-walled structures like arteries. For example, [22] introduced a layer-specific axisymmetric shell formulation within a constrained mixture model to study arterial growth and remodeling. [23] used a Kirchhoff–Love shell in an isogeometric framework for cardiac electromechanics, leveraging assumptions of inextensible directors and negligible transverse shear. [24, 25] developed a nonlinear, rotation-free shell formulation based on the Kirchhoff–Love theory, utilizing midsurface displacement and a director vector to capture thickness stretch, which is statically condensed to enforce incompressibility via a plane stress condition. [25] also compared their shell formulation with membrane and solid elements in vascular bifurcations, demonstrating that the shell approach offers an optimal balance between computational cost and accuracy. Nonetheless, these vascular and cardiac findings do not directly transfer to gastric motility. The stomach wall is thicker than vascular tissue and undergoes significant local contractions with notable thickness changes during peristalsis. Cardiac muscle, while thick, typically deforms more uniformly and it thus more suitable for traditional shell assumptions.
In summary, the persisting challenges in whole-organ coupled electromechanical modeling of the human stomach include the treatment of large nonlinear deformations, the treatment of the multiple spatially heterogeneous tissue properties, and the accommodation of substantial inter-individual variability. These variabilities encompass patient-specific geometries and complex, heterogeneous spatial propagation and motility patterns [26].
@@
As in most biomechanical applications, the intrinsic multiscale nature of the system requires ad hoc simplifying assumptions, which allow the construction of a reliable computational framework to advance understanding of a specific subproblem. Accordingly, in this paper, we focus on the electromechanical modeling and computational implementation of gastric motility, while disregarding fluid-structure interaction. Building on our previous work [16], we adopt a phenomenological approach and generalize it within the framework of constrained mixture theory [27], incorporating an anisotropic active strain formulation [28, 29]. To address the challenges posed by large deformations and nonlinear material behavior in the gastric wall, we transition from the previously employed membrane-based formulation to a $7$-parameter shell model [30, 31], which captures thickness stretch and transverse shear. Additionally, we introduce a whole-organ gastric motility model that accounts for heterogeneous parameter fields. These fields encode regional variations in key physiological features, including intrinsic slow wave frequency, electrical conductivity, tissue stiffness, and active contractility, across the full gastric domain. Unlike previous approaches that rely on discrete anatomical subregions or piecewise-constant parameters, our framework defines smooth parameter gradients informed by anatomy and physiology. This enables realistic simulation of entrainment, wave propagation, and peristaltic contractions, while preserving continuity and anatomical fidelity throughout the domain.
We demonstrate the capabilities of our model by simulating a realistic human stomach geometry, incorporating prestress, regionally varying material properties, and non-uniform boundary conditions. The study highlights the improved performance of the $7$-parameter shell compared to membrane and solid element formulations through a rigorous computational comparison of electromechanical motility and demonstrates how regionally varying heterogeneous fields enable spatially accurate simulation of contraction patterns. We provide an open-source implementation of our formulation within the multiphysics simulation framework 4C [32].
@@ 2 Electromechanical model
We start by setting the mathematical and mechanical framework which is based on the general theory of nonlinear continuum mechanics (e.g. [33]). A material point $\boldsymbol{X}$ in the reference configuration $\Omega_{\textup{R}}$ is mapped to a spatial position $\boldsymbol{x}$ in the current configuration $\Omega_{\textup{t}}$ at time $t$ via
@@ 2.1 Structural model: nonlinear 7-parameter shell
The stomach is a thin-walled, curved structure (see Figure 1), which aligns with the core assumption of shell formulations: one dimension (the thickness) is much smaller than the other two (in-plane) dimensions. Based on this assumption, we briefly summarize the mathematical formulation of the $7$-parameter shell originally proposed in [30, 31].
@@ 2.1.2 Kinematics
According to the enhanced assumed strain (EAS) method [34, 35], an independent strain field $\boldsymbol{\tilde{E}}$ is introduced to avoid artificial stresses in the thickness direction, resulting in the total Green-Lagrange strain tensor $\boldsymbol{E}=\boldsymbol{E}^{u}+\boldsymbol{\tilde{E}}$. The additional strains affect only the transversal normal strains, i.e.,
@@ 2.1.3 Constitutive model
where $\hat{\mu}$ represents the determinant of the shell shifter tensor [30, 31], which transforms a volume integral into an integration over the shell midsurface.
Leveraging these definitions and employing the principle of internal virtual work [33], the static mechanical equilibrium of the shell, considering the differential shell volume $\mathrm{d}V=\hat{\mu}\ \mathrm{d}\theta^{3}\mathrm{d}A$, is expressed as
@@ 2.2.1 General
The gastric microstructure is represented as a mixture of three load-bearing constituents: an isotropic ground matrix, mainly elastin, and two fiber families composed of collagen and SMCs. The fiber families are oriented in longitudinal and circumferential direction respectively, cf. Figures 4d and 4e. For simplicity, we do not distinguish between collagen and smooth muscle fibers. Instead, we assume that the fiber families in our model represent both. Each constituent is modeled individually and characterized by its mass fraction, material properties, and the prestress it exhibits in the in vivo reference configuration [27]. We denote quantities associated with a specific constituent by a superscript $i\in I$, where $I$ is an index set of cardinality $N$, corresponding to the number of constituents.
@@
Here, $\boldsymbol{F}_{\textup{e}}^{i}$ represents the elastic part of the deformation. $\boldsymbol{F}_{\textup{a}}^{i}$ models that gastric tissue exhibits active muscular contractions. The resulting deformation is not an elastic one but an active (inelastic) change of the natural configuration of a tissue volume element. It can be accounted for by a so-called active strain approach [28, 29] through an active part $\boldsymbol{F}_{\textup{a}}^{i}$ of the deformation gradient. Ultimately, one has to account for the fact that biological tissues are in general able to grow and remodel. Growth and remodeling induces permanent, inelastic changes of the stress-free natural configuration of a tissue and can be modeled by a related inelastic part of the deformation gradient $\boldsymbol{F}_{\textup{gr}}^{i}$ [36]. In the following, we describe in detail how the individual components of the deformation gradient are defined and computed.
@@ 2.2.2 Reference and initial configuration
Here, $\boldsymbol{F}(0)$ and $\boldsymbol{F}_{\textup{e}}^{i}(0)$ denote the deformation gradient at $t=0$ and its elastic part, respectively. The requirement $\boldsymbol{F}(0)=\boldsymbol{I}$ results from our definition that in $\Omega_{\textup{R}}$ and $\Omega_{0}$ our body exhibits the same geometry. That is, in the fictitious configuration $\Omega_{\textup{R}}$, each material point lies at the same position as in the initial configuration $\Omega_{0}$ acquired from medical imaging. Herein, we do not study growth and remodeling of gastric tissue. That is, we assume that such processes do not happen on the time scale of peristaltic contractions that is relevant for this article. This is motivated by the fact that growth and remodeling typically happens on the time scale of weeks to months, whereas peristaltic contractions happen on the time scale of minutes to hours. Therefore, the inelastic part of the deformation gradient due to growth and remodeling, $\boldsymbol{F}_{\textup{gr}}^{i}$, captures only effects of previous growth and remodeling in the past. It can thus be assumed in the above Eq. 15 to be a constant in time. To determine this constant for each constituent, we essentially rely on concepts for prestress computation proposed by [37] and [38] and recently used also by [39]. Details are discussed below.
@@ 2.2.3 Material elasticity
In our stomach model, anisotropic stiffness contributions arise from smooth muscle and collagen fiber families arranged in preferential directions (longitudinal and circumferential), whereas the isotropic ground matrix is attributed to the elastin network. For sake of conciseness, the specific form of the strain energy functions $W^{i}$ used to characterize each constituent are provided in A. These expressions follow established formulations (e.g. [40, 33]) and are not repeated here to focus on model integration and physiological specificity.

### v2

@@ Abstract
[5]organization= Theoretical and Computational Biomechanics Lab, Department of Engineering, Università Campus Bio-Medico di Roma, addressline=Via A. del Portillo 21, city=Rome, postcode=00128, country=Italy
@@ 1 Introduction
The stomach plays a central role in the gastrointestinal tract. It is concerned with storing, mixing, and grinding food. Fundamental to this functionality are rhythmic gastric contractions, which generate peristaltic waves that mix and propel food through the digestive tract. Perturbations of this system are associated with a broad spectrum of diseases and disorders, including obesity, which itself is a major risk factor for conditions such as cardiovascular disease. Other gastric motility disorders include gastroparesis, gastroesophageal reflux disease (GERD), and functional dyspepsia, all of which are linked to abnormal gastric motility [Streutker-2007-ICCHealthDisease, Grady-2021-GastricConductionReview]. A detailed understanding of gastric motility is therefore essential for the effective diagnosis and treatment.
@@
Anatomically, the stomach is divided into three main regions (fundus, corpus, antrum) and exhibits both a so-called greater and lesser curvature as illustrated in Figure 1. Despite its relatively thin wall (on average $2.84.5\text{\,}\frac{\mathrm{mm}}{}$ in humans [Friis-2023-GastricTissue, Holzer-Stock2025]) compared to its overall size (approximately $1525\text{\,}\frac{\mathrm{cm}}{}$ in length [Friis-2023-GastricTissue, Holzer-Stock2025]), the gastric wall consists of distinct layers. The muscular layer, primarily responsible for gastric contractions, comprises two smooth muscle sheets: one oriented circumferentially and the other longitudinally. An additional oblique smooth muscle layer is restricted to the upper corpus near the esophagus. It is neglected in this study, as it plays a limited role in gastric peristalsis, which occurs mainly in the lower body and antrum [DiNatale-2023-FunctionalAndAnatomicalRegions].
@@
The stomach exhibits intricate patterns of mechanical contraction [DiNatale-2023-FunctionalAndAnatomicalRegions]. Gastric motility is governed by complex electromechanical couplings comprising at least two cell types, interstitial cells of Cajal (ICCs) and smooth muscle cells (SMCs) [huizinga2009a]. ICCs, arranged in a distributed network, act as pacemaker cells generating electrical activity in the form of slow waves [Sanders-2016-RegulationSMCFunction]. Being self-excitatory, ICCs show peculiar features of nonlinear coupled oscillators [Winfree1987, Glass2001]; i.e., they generate electrical activity at specific intrinsic frequencies while the integrated network synchronizes at a unified emerging frequency. This so-called entrainment generates ordered waves at the highest intrinsic frequency of approximately $3\text{\,}\frac{\mathrm{cpm}}{}$ (cycles per minute). The highest intrinsic frequency is observed near the pacemaker region at the mid-corpus along the greater curvature. The intrinsic frequency decreases longitudinally and circumferentially and assumes its minimum near the pylorus at the lesser curvature [Grady-2021-GastricConductionReview]. The SMCs are electrically coupled to ICCs, such that the electrical slow waves coordinate the contractions of the SMCs via an elaborate excitation–contraction mechanism [CorriasBuist-2007-SMCActivation, sanders2012b]. These contractions can produce substantial gastric wall deformations, with amplitudes measured in the range of $720\text{\,}\frac{\mathrm{mm}}{}$ [Schulze-2006-ImagingStomach, Hosseini2023, Wang2024].
@@
Early whole-organ modeling approaches of the stomach focused primarily on its electrophysiological system [du2013c]. More recently, models incorporating mechanical phenomena have also been proposed. Our previous work [brandstaeter2018a] was the first to simulate gastric electromechanics on an idealized whole-organ geometry, reproducing key phenomena. However, it relied on a highly simplified geometry and material model of the tissue and could deal only with very limited contraction amplitudes due to reduced numerical robustness that resulted from the use of membrane finite elements to discretize the thin wall of the stomach. Subsequent developments have advanced mechanical modeling toward anatomically realistic stomach geometries. [Fontanella2019] examined gastric deformation in animal-specific models but were restricted to passive tissue mechanics. Similarly, [Toniolo2022a, Toniolo2024] applied a passive mechanical framework to analyze patient-specific bariatric surgery, precluding the simulation of active motility. In contrast, [Klemm2020a, Klemm2023] proposed an electro-chemomechanical model for the porcine stomach, using an active-stress formulation and incorporating mechano-electric feedback to capture propagating contraction waves. While their approach was particularly advanced in terms of the electro-chemical components, several limitations remained such as an unrealistic antrum activation and mild contraction intensity.
@@
The thin-walled structure of the stomach lends itself naturally to reduced-dimensional formulations, such as membrane or shell models. While membrane models are computationally efficient, they neglect bending stiffness, limiting their application in regions with high curvature and large contractions. Nonlinear shell formulations capture both membrane and bending effects under large deformations and are thus preferable for accurate gastric biomechanics. Shell models have proven effective for studying the biomechanics of thin-walled structures like arteries. For example, [Laubrie2020] introduced a layer-specific axisymmetric shell formulation within a constrained mixture model to study arterial growth and remodeling. [Nitti2021] used a Kirchhoff–Love shell in an isogeometric framework for cardiac atrial electromechanics, leveraging assumptions of inextensible directors and negligible transverse shear. [Nama2020, Nama2023] developed a nonlinear, rotation-free shell formulation for vascular biomechanics based on the Kirchhoff–Love theory, utilizing midsurface displacement and a director vector to capture thickness stretch, which is statically condensed to enforce incompressibility via a plane stress condition. [Nama2023] also compared their shell formulation with membrane and solid elements in vascular bifurcations, demonstrating that the shell approach offers an optimal balance between computational cost and accuracy.
However, these cardiovascular findings cannot be directly translated to gastric tissue. Although the stomach wall is thicker than vascular tissues—where shell models have been successfully applied—it remains thinner than the ventricular wall of the heart, which typically requires solid finite element formulations. Cardiac ventricular models must capture thickness- and sheet-dependent features such as fiber rotational anisotropy, necessitating advanced, structure-based modeling frameworks [Holzapfel-2009-ConstitutiveMyocard, Eriksson2013, Sommer2015, Gltekin2016]. The substantial wall thickness of the cardiac chambers, particularly the left ventricle, demands solid finite element to achieve physiologically accurate and mechanically reliable simulations [Land2015, Viola2020, Brown2025, Aróstica-2025-CardiacBenchmark].
In summary, the persisting challenges in whole-organ coupled electromechanical modeling of the human stomach include the treatment of large nonlinear deformations, the treatment of the multiple spatially heterogeneous tissue properties, and the accommodation of substantial inter-individual variability. These variabilities encompass patient-specific geometries and complex, heterogeneous spatial propagation and motility patterns [EgbujiGrady-2010-OriginPropagationSW].
@@
As in most biomechanical applications, the intrinsic multiscale nature of the system requires ad hoc simplifying assumptions, which allow the construction of a reliable computational framework to advance understanding of a specific subproblem. Accordingly, in this paper, we focus on the electromechanical modeling and computational implementation of gastric motility, while disregarding fluid-structure interaction. Building on our previous work [brandstaeter2018a], which introduced a preliminary electromechanical framework limited to basic material models, simplified geometries, and reduced contractions, we substantially extend the approach in the present study. Here, we generalize the formulation within the framework of constrained mixture theory [Humphrey-2021-CMMOverview], incorporate an anisotropic active strain model [Brandstaeter2019, Patel-Gizzi-2022-GastrointestinalTissueReview], and integrate spatially varying electrophysiological and mechanical fields derived automatically from the organ geometry. To address the challenges posed by large deformations and nonlinear material behavior in the gastric wall, we transition from the previously employed membrane-based formulation to a $7$-parameter shell model [Braun-1994-Shell, Büchter-1994-3DExtensionEAS], which captures thickness stretch and transverse shear. Additionally, we introduce a whole-organ gastric motility model that accounts for heterogeneous parameter fields. These fields encode regional variations in key physiological features, including intrinsic slow wave frequency, electrical conductivity, tissue stiffness, and active contractility, across the full gastric domain. Unlike previous approaches that rely on discrete anatomical subregions or piecewise-constant parameters, our framework defines smooth parameter gradients informed by anatomy and physiology. This enables realistic simulation of entrainment, wave propagation, and peristaltic contractions, while preserving continuity and anatomical fidelity throughout the domain.
We demonstrate the capabilities of our model by simulating a realistic human stomach geometry, incorporating prestress, regionally varying material properties, and non-uniform boundary conditions. The study highlights the improved performance of the $7$-parameter shell compared to membrane and solid element formulations through a rigorous computational comparison of electromechanical motility and demonstrates how regionally varying heterogeneous fields enable spatially accurate simulation of contraction patterns. We provide an open-source implementation of our formulation within the multiphysics simulation framework 4C [4C].
@@ 2 Electromechanical model
We start by setting the mathematical and mechanical framework which is based on the general theory of nonlinear continuum mechanics (e.g. [Holzapfel-2001-NonlinearSolidMechanics]). A material point $\boldsymbol{X}$ in the reference configuration $\Omega_{\textup{R}}$ is mapped to a spatial position $\boldsymbol{x}$ in the current configuration $\Omega_{\textup{t}}$ at time $t$ via
@@ 2.1 Structural model: nonlinear 7-parameter shell
The stomach is a thin-walled, curved structure (see Figure 1), which aligns with the core assumption of shell formulations: one dimension (the thickness) is much smaller than the other two (in-plane) dimensions. Based on this assumption, we briefly summarize the mathematical formulation of the $7$-parameter shell originally proposed in [Braun-1994-Shell, Büchter-1994-3DExtensionEAS].
@@ 2.1.2 Kinematics
According to the enhanced assumed strain (EAS) method [SimoRifai-1990-EAS, Büchter-1994-3DExtensionEAS], an independent strain field $\boldsymbol{\tilde{E}}$ is introduced to avoid artificial stresses in the thickness direction, resulting in the total Green-Lagrange strain tensor $\boldsymbol{E}=\boldsymbol{E}^{u}+\boldsymbol{\tilde{E}}$. The additional strains affect only the transversal normal strains, i.e.,
@@ 2.1.3 Constitutive model
where $\hat{\mu}$ represents the determinant of the shell shifter tensor [Braun-1994-Shell, Büchter-1994-3DExtensionEAS], which transforms a volume integral into an integration over the shell midsurface.
Leveraging these definitions and employing the principle of internal virtual work [Holzapfel-2001-NonlinearSolidMechanics], the static mechanical equilibrium of the shell, considering the differential shell volume $\mathrm{d}V=\hat{\mu}\ \mathrm{d}\theta^{3}\mathrm{d}A$, is expressed as
@@ 2.2.1 General
The gastric microstructure is represented as a mixture of three load-bearing constituents: an isotropic ground matrix, mainly elastin, and two fiber families composed of collagen and SMCs. The fiber families are oriented in longitudinal and circumferential direction respectively, cf. Figures 4d and 4e. For simplicity, we do not distinguish between collagen and smooth muscle fibers. Instead, we assume that the fiber families in our model represent both. Each constituent is modeled individually and characterized by its mass fraction, material properties, and the prestress it exhibits in the in vivo reference configuration [Humphrey-2021-CMMOverview]. We denote quantities associated with a specific constituent by a superscript $i\in I$, where $I$ is an index set of cardinality $N$, corresponding to the number of constituents.
@@
Here, $\boldsymbol{F}_{\textup{e}}^{i}$ represents the elastic part of the deformation. $\boldsymbol{F}_{\textup{a}}^{i}$ models that gastric tissue exhibits active muscular contractions. The resulting deformation is not an elastic one but an active (inelastic) change of the natural configuration of a tissue volume element. It can be accounted for by a so-called active strain approach [Brandstaeter2019, Patel-Gizzi-2022-GastrointestinalTissueReview] through an active part $\boldsymbol{F}_{\textup{a}}^{i}$ of the deformation gradient. Ultimately, one has to account for the fact that biological tissues are in general able to grow and remodel. Growth and remodeling induces permanent, inelastic changes of the stress-free natural configuration of a tissue and can be modeled by a related inelastic part of the deformation gradient $\boldsymbol{F}_{\textup{gr}}^{i}$ [Braeu-2017-HCMM]. In the following, we describe in detail how the individual components of the deformation gradient are defined and computed.
@@ 2.2.2 Reference and initial configuration
Here, $\boldsymbol{F}(0)$ and $\boldsymbol{F}_{\textup{e}}^{i}(0)$ denote the deformation gradient at $t=0$ and its elastic part, respectively. The requirement $\boldsymbol{F}(0)=\boldsymbol{I}$ results from our definition that in $\Omega_{\textup{R}}$ and $\Omega_{0}$ our body exhibits the same geometry. That is, in the fictitious configuration $\Omega_{\textup{R}}$, each material point lies at the same position as in the initial configuration $\Omega_{0}$ acquired from medical imaging. Herein, we do not study growth and remodeling of gastric tissue. That is, we assume that such processes do not happen on the time scale of peristaltic contractions that is relevant for this article. This is motivated by the fact that growth and remodeling typically happens on the time scale of weeks to months, whereas peristaltic contractions happen on the time scale of minutes to hours. Therefore, the inelastic part of the deformation gradient due to growth and remodeling, $\boldsymbol{F}_{\textup{gr}}^{i}$, captures only effects of previous growth and remodeling in the past. It can thus be assumed in the above Eq. 15 to be a constant in time. To determine this constant for each constituent, we essentially rely on concepts for prestress computation proposed by [Weisbecker-2014-Prestress] and [MousaviAvril-2017] and recently used also by [Gebauer-2023-CMMCardiac]. Details are discussed below.
@@ 2.2.3 Material elasticity
In our stomach model, anisotropic stiffness contributions arise from smooth muscle and collagen fiber families arranged in preferential directions (longitudinal and circumferential), whereas the isotropic ground matrix is attributed to the elastin network. For sake of conciseness, the specific form of the strain energy functions $W^{i}$ used to characterize each constituent are provided in A. These expressions follow established formulations (e.g. [HolzapfelGasserOgden-2012-ConstArterial, Holzapfel-2001-NonlinearSolidMechanics]) and are not repeated here to focus on model integration and physiological specificity.